Leapin’ Larry’s Pre-Owned Cars has two divisions, Operations and Financing. Operations is responsible for selling Larry’s inventory as quickly as possible and purchasing cars for future sale. Financing Division takes loan applications and packages loans into pools and sells them in the financial markets. It also services the loans. Both divisions meet the requirements for segment disclosures under accounting rules.

 

Operations Division had $62 million in sales last year. Costs, other than those charged by Financing Division, totaled $28 million. Financing Division earned revenues of $19 million from servicing loans and incurred outside costs of $20 million. In addition, Financing charged Operations $18 million for loan-related fees. Operations’s manager complained to Larry that Financing was charging twice the commercial rate for loan-related fees and that Operations would be better off sending its buyers to an outside lender.

 

Financing's manager replied that although commercial rates could be lower, servicing Larry's loans is more difficult, thereby justifying the higher fees.

 

Required:

a. What are the reported segment operating profits for each division, ignoring income taxes and using the $18 million transfer price for the loan-related fees?

b. What are the reported segment operating profits for each division, ignoring income taxes and using a $9 million commercial rate as the transfer price for the loan-related fees?
Answer #1

3A) 62-28-18= 16 operating division

19-20+18= 17 financing division 


3B) 62-28-9= 25 operating division

19-20+9= 8 financing division